OpenStudy (anonymous):

3 times the measure of an angle is 14 less than the measure of its complement. What is the measure of the angle? A.19° B.57° C.71° D.76°

OpenStudy (anonymous):

19. 19*3=57 the complement for 19 is 71 71-57=14. :)

OpenStudy (anonymous):

3A = 90 - A - 19 -> A = 76/4 = ...